### v1

# tablesection algorithmsection

Abstract
This paper studies two estimators for Gaussian moment tensors: the standard sample moment estimator and a plug-in estimator based on Isserlis’s theorem. We establish dimension-free, non-asymptotic error bounds that demonstrate and quantify the advantage of Isserlis’s estimator for tensors of even order $p>2.$ Our bounds hold in operator and entrywise maximum norms, and apply to symmetric and asymmetric tensors.
11footnotetext: Eric and Wendy Schmidt Center, Broad Institute of MIT and Harvard22footnotetext: Committee on Computational and Applied Mathematics, University of Chicago33footnotetext: Department of Statistics, University of Chicago
1 Introduction
This paper develops dimension-free, non-asymptotic theory for estimating Gaussian moment tensors of arbitrary order. We study two estimators: the sample moment estimator and a plug-in estimator based on Isserlis’s formula. Our non-asymptotic bounds quantify the advantage of Isserlis’s estimator for tensors of even order $p>2,$ both in operator and entrywise maximum norms. By studying the sample moment estimator, this work contributes to the emerging literature on concentration inequalities for random tensors [27, 30, 34, 33, 9, 11, 4, 12]; and by studying Isserlis’s estimator, it contributes to the literature on estimating functionals of high and infinite-dimensional parameters [16, 17, 20, 18]. More broadly, this work is motivated by the growing adoption of tensor methods in statistics and machine learning [23, 21, 8, 10, 7, 13, 25], which calls for new theory for high-dimensional tensor estimation.
The concentration of the sample covariance (the sample moment tensor of order $p=2$) has been extensively studied. Most pertinent to our work, the papers [28, 19] identify a notion of effective dimension that determines the sample complexity. In contrast to our results for even order $p>2,$ in the case $p=2$ the sample covariance matrix is known to be minimax optimal under the operator norm (see [22, Theorem 2] and [19, Theorem 4]). Sharp operator-norm concentration inequalities for simple random tensors of arbitrary order were recently established in [2, 12] using the effective dimension from [28, 19]. In this paper, we prove a new sharp concentration inequality in entrywise maximum norm using a notion of effective dimension introduced in [3, 4] to analyze thresholding estimators for structured covariance operators.
Isserlis’s theorem expresses higher-order moments of a Gaussian distribution as functionals of its covariance. Consequently, our problem falls within the general framework of estimating functionals $f(\Sigma)$ of a covariance operator $\Sigma$ [16, 17], where plug-in estimators $f(\widehat{\Sigma})$ based on the sample covariance $\widehat{\Sigma}$ are often severely biased and suboptimal [20, 18]. Bias reduction techniques —such as jackknife, bootstrap, and Taylor expansions— are commonly used to improve efficiency. Our new dimension-free, non-asymptotic error bounds for Isserlis’s estimator show that, in the context of estimating Gaussian moment tensors, plug-in estimation via Isserlis’s formula performs well and strictly outperforms the sample moment estimator.
2 Preliminaries
Given positive integers $d_{1},\dots,d_{p}$ and tensors $T,T^{\prime}\in\mathbb{R}^{d_{1}\times\cdots\times d_{p}}$, their Frobenius inner product is defined by
$\mathopen{}\mathclose{{}\left\langle T,T^{\prime}}\right\rangle:=\sum_{i_{1}=1}^{d_{1}}\sum_{i_{2}=1}^{d_{2}}\cdots\sum_{i_{p}=1}^{d_{p}}T_{i_{1},\dots,i_{p}}T^{\prime}_{i_{1},\dots,i_{p}}.$
For vectors $v_{k}\in\mathbb{R}^{d_{k}},$ $1\leq k\leq p,$ let $v_{1}\otimes\cdots\otimes v_{p}\in\mathbb{R}^{d_{1}\times\cdots\times d_{p}}$ denote their outer product, defined entrywise by
$(v_{1}\otimes\cdots\otimes v_{p})_{\ell_{1},\dots,\ell_{p}}:=v_{1,\ell_{1}}\cdots v_{p,\ell_{p}}.$
The corresponding operator norm is given by
$\|T\|:=\sup_{v_{k}\in\mathcal{S}^{d_{k}},1\leq k\leq p}\mathopen{}\mathclose{{}\left\langle T,v_{1}\otimes\cdots\otimes v_{p}}\right\rangle,$
where $\mathcal{S}^{d_{k}}:=\{v\in\mathbb{R}^{d_{k}}:|v|=1\}$ denotes the unit Euclidean sphere in $\mathbb{R}^{d_{k}}.$ Similarly, the entrywise maximum norm is defined by
$\|T\|_{\max}:=\sup_{v_{k}\in\mathcal{E}_{k},1\leq k\leq p}|\mathopen{}\mathclose{{}\left\langle T,v_{1}\otimes\cdots\otimes v_{p}}\right\rangle|,$
where $\mathcal{E}_{k}:=\{e_{i}\}_{i=1}^{d_{k}}$ is the standard basis of $\mathbb{R}^{d_{k}}.$
We will consider two notions of effective dimension of a covariance matrix $\Sigma\in\mathbb{R}^{d\times d}.$ The first is
$r_{2}(\Sigma):=\frac{\text{Tr}(\Sigma)}{\|\Sigma\|},$
which will arise in our bounds in operator norm (see e.g. [19, 28]). The second is
$r_{\max}(\Sigma):=\frac{\mathopen{}\mathclose{{}\left(\mathbb{E}_{X\sim\mathcal{N}(0,\Sigma)}\|X\|_{\infty}}\right)^{2}}{\|\Sigma\|_{\max}},$
which arises in our bounds in entrywise maximum norm (see e.g. [5, 3, 4]).
For positive sequences $\{a_{n}\},\{b_{n}\}$, we write $a_{n}\lesssim b_{n}$ to denote that, for some constant $c>0$, $a_{n}\leq cb_{n}.$ If both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold, we write $a_{n}\asymp b_{n}.$ We write $\lesssim_{p}$ and $\asymp_{p}$ to indicate that the implicit constant may depend on the parameter $p$.
3 Main results
This section contains our main results. In Subsection 3.1, we study the symmetric case: moment tensors defined by expectation of the $p$-fold outer product of a single random vector. In Subsection 3.2, we consider the asymmetric case: moment tensors defined by expectation of the outer product of $p$ random vectors of possibly different dimensions. While the asymmetric case subsumes the symmetric one, we present first the symmetric case due to its simplicity and its central role in theory and applications.
3.1 Symmetric case
Let $X\sim\mathcal{N}(0,\Sigma)$ be a zero-mean Gaussian random vector in $\mathbb{R}^{d}$ with covariance matrix $\Sigma,$ and let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$. Let $p$ be an even integer. Our goal is to estimate the $p$-th order moment tensor
$T:=\mathbb{E}X^{\otimes p}:=\mathbb{E}X\otimes\cdots\otimes X,$
which is the expectation of the $p$-fold tensor product of $X$.
A natural estimator is the sample moment estimator $\widehat{T}_{S}$, defined by
$\widehat{T}_{S}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}.$
Entrywise, this corresponds to
$\big{(}\widehat{T}_{S}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\frac{1}{N}\sum_{i=1}^{N}X_{i,\ell_{1}}\cdots X_{i,\ell_{p}},$
where $X_{i,\ell}$ denotes the $\ell$-th coordinate of $X_{i}.$ Alternatively, Isserlis’s theorem [15] (also known as Wick’s probability theorem [31]) provides an exact expression for Gaussian moment tensors in terms of second-order moments. Specifically, for any multi-index $(\ell_{1},\dots,\ell_{p}),$
$(\mathbb{E}X^{\otimes p})_{\ell_{1},\dots,\ell_{p}}=\mathbb{E}[X_{\ell_{1}}\cdots X_{\ell_{p}}]=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\Sigma_{\ell_{j},\ell_{k}},$
where $\Pi^{2}_{p}$ is the set of all pairwise partitions of $\{1,\dots,p\}$. This motivates Isserlis’s estimator, which substitutes the sample covariance
$\widehat{\Sigma}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}X_{i}^{\top}$
into the same expression. The resulting estimator is given entrywise by
$\big{(}\widehat{T}_{I}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\widehat{\Sigma}_{\ell_{j},\ell_{k}}.$
Using the notion of induced likelihood [32], the plug-in estimator $\widehat{T}_{I}$ can be interpreted as the maximum likelihood estimator of $T.$
Theorem 3.1 compares the performance of the two estimators, $\widehat{T}_{S}$ and $\widehat{T}_{I}$, under both the operator norm and the entrywise maximum norm.
Theorem 3.1.
Using the notation above, the following bounds hold:
(i) Operator norm bounds:
$\mathbb{E}\lVert\widehat{T}_{S}-T\rVert$ $\asymp_{p}\|\Sigma\|^{p/2}\bigg{(}\sqrt{\frac{r_{2}(\Sigma)}{N}}+\frac{r_{2}(\Sigma)^{p/2}}{N}\bigg{)},$
$\mathbb{E}\lVert\widehat{T}_{I}-T\rVert$ $\lesssim_{p}\|\Sigma\|^{p/2}\bigg{(}\sqrt{\frac{r_{2}(\Sigma)}{N}}+\bigg{(}\frac{r_{2}(\Sigma)}{N}\bigg{)}^{p/2}\bigg{)}.$
(ii) Entrywise maximum norm bounds:
$\mathbb{E}\lVert\widehat{T}_{S}-T\rVert_{\max}$ $\asymp_{p}\|\Sigma\|_{\max}^{p/2}\bigg{(}\sqrt{\frac{r_{\max}(\Sigma)}{N}}+\frac{r_{\max}(\Sigma)^{p/2}}{N}\bigg{)},$
$\mathbb{E}\lVert\widehat{T}_{I}-T\rVert_{\max}$ $\lesssim_{p}\|\Sigma\|_{\max}^{p/2}\bigg{(}\sqrt{\frac{r_{\max}(\Sigma)}{N}}+\bigg{(}\frac{r_{\max}(\Sigma)}{N}\bigg{)}^{p/2}\bigg{)}.$
Proof.
(i) Operator norm bounds. The sample moment estimator bound follows by [2, Theorem 2.1] (see also [1]). For $\widehat{T}_{I}$, by Corollary 3.7, we have
$\frac{\|\widehat{T}_{I}-T\|}{\|T\|}\leq\frac{p}{2}\cdot\frac{\|\widehat{\Sigma}-\Sigma\|}{\|\Sigma\|}\bigg{(}1+\frac{\|\widehat{\Sigma}-\Sigma\|}{\|\Sigma\|}\bigg{)}^{p/2-1}\lesssim_{p}\frac{\|\widehat{\Sigma}-\Sigma\|}{\|\Sigma\|}+\bigg{(}\frac{\|\widehat{\Sigma}-\Sigma\|}{\|\Sigma\|}\bigg{)}^{p/2}.$
Further, by [19, Theorem 4], $\mathbb{E}\|\widehat{\Sigma}-\Sigma\|/\|\Sigma\|\asymp\sqrt{\frac{r_{2}(\Sigma)}{N}}+\frac{r_{2}(\Sigma)}{N},$ which can be plugged into the expression in the previous display to give the result.
(ii) Entrywise maximum norm bounds. For the sample moment estimator, the upper bound is a direct corollary of Theorem 3.4. The lower bound follows by an analogous argument to the one used in the proof of [2, Theorem 2.1] and is omitted for brevity. For Isserlis’s estimator, the bound follows from Corollary 3.7. ∎
Remark 3.2.
Theorem 3.1 shows that consistent estimation of $T$ using the sample moment estimator $\widehat{T}_{S}$ requires a sample size $N\gg(r_{2}(\Sigma))^{p/2}$ under the operator norm, or $N\gg(r_{\max}(\Sigma))^{p/2}$ under the entrywise maximum norm. In contrast, consistency of Isserlis’s estimator $\widehat{T}_{I}$ only requires sample size $N\gg r_{2}(\Sigma)$ or $N\gg r_{\max}(\Sigma)$, leading to a significant reduction in sample complexity. ∎
Remark 3.3.
[24, Theorem 3] generalizes Isserlis’s theorem to isotropic distributions. While beyond the scope of this work, we anticipate the advantage of Isserlis’s estimators to extend to that setting. Numerical results in [6] support this claim. ∎
3.2 Asymmetric case
Let $p$ be an even integer, and let $X=(X^{(1)},\ldots,X^{(p)})\in\mathbb{R}^{d}$ be zero-mean Gaussian random vectors with covariance matrix $\Sigma$, where each block $X^{(k)}\in\mathbb{R}^{d_{k}}$ and $\sum_{k=1}^{p}d_{k}=:d$. For each $k\in\{1,\ldots,p\}$, denote the marginal covariance by $\Sigma^{(k)}:=\mathbb{E}X^{(k)}\otimes X^{(k)}$, and the cross-covariance by $\Sigma^{(j,k)}:=\mathbb{E}X^{(j)}\otimes X^{(k)}$. Let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$, where $X_{i}=(X^{(1)}_{i},\ldots,X^{(p)}_{i})$. Our goal is to estimate the moment tensor
$T:=\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)},$
formed by taking the tensor product over the blocks of $X$.
A natural estimator is the sample moment estimator, defined as
$\widehat{T}_{S}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}.$
Entrywise, this corresponds to
$\big{(}\widehat{T}_{S}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\frac{1}{N}\sum_{i=1}^{N}X^{(1)}_{i,\ell_{1}}\cdots X^{(p)}_{i,\ell_{p}},$
where $X^{(k)}_{i,\ell}$ denotes the $\ell$-th coordinate of $X^{(k)}_{i}.$
Alternatively, by Isserlis’s theorem, for any multi-index $(\ell_{1},\ldots,\ell_{p})$,
$\big{(}\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}\big{)}_{\ell_{1},\dots,\ell_{p}}=\mathbb{E}[X^{(1)}_{\ell_{1}}\cdots X^{(p)}_{\ell_{p}}]=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\Sigma^{(j,k)}_{\ell_{j},\ell_{k}},$
where $\Pi^{2}_{p}$ denotes the set of pairwise partitions of $\{1,\ldots,p\}$. This motivates Isserlis’s estimator, which substitutes the sample covariances
$\widehat{\Sigma}^{(j,k)}:=\frac{1}{N}\sum_{i=1}^{N}X^{(j)}_{i}(X^{(k)}_{i})^{\top}$
into the same expression. The resulting estimator is given entrywise by
$\big{(}\widehat{T}_{I}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\widehat{\Sigma}^{(j,k)}_{\ell_{j},\ell_{k}}.$
Theorem 3.4 compares the performance of the two estimators, $\widehat{T}_{S}$ and $\widehat{T}_{I}$, under both the operator norm and the entrywise maximum norm.
Theorem 3.4.
Using the notation above, the following bounds hold:
(i) Operator norm bounds:
$\mathbb{E}\|\widehat{T}_{S}-T\|$ $\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}\Bigg{(}\bigg{(}\frac{\sum_{k=1}^{p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+\frac{\prod_{k=1}^{p}(r_{2}(\Sigma^{(k)})+\log N)^{1/2}}{N}\Bigg{)},$
$\mathbb{E}\|\widehat{T}_{I}-T\|$ $\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2},\ \ \text{if }\ N\geq\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)}).$
(ii) Entrywise maximum norm bounds:
$\mathbb{E}\|\widehat{T}_{S}-T\|_{\max}$ $\lesssim_{p}\!\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|_{\max}^{1/2}\bigg{)}\Bigg{(}\bigg{(}\frac{\sum_{k=1}^{p}r_{\max}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}\!\!+\frac{\prod_{k=1}^{p}(r_{\max}(\Sigma^{(k)})+\log N)^{1/2}}{N}\Bigg{)},$
$\mathbb{E}\|\widehat{T}_{I}-T\|_{\max}$ $\lesssim_{p}\!\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|_{\max}^{1/2}\bigg{)}\bigg{(}\frac{\max_{1\leq k\leq p}r_{\max}(\Sigma^{(k)})}{N}\bigg{)}^{1/2},\ \ \text{if }\ N\geq\max_{1\leq k\leq p}r_{\max}(\Sigma^{(k)}).$
Remark 3.5.
The upper bounds on the deviation of sample moment estimator $\widehat{T}_{S}$ from its expectation are sharp under both the operator norm and entrywise maximum norm when the vectors $X^{(1)},\ldots,X^{(p)}$, together with their samples $(X^{(1)}_{i})_{i=1}^{N},\ldots,(X^{(p)}_{i})_{i=1}^{N}$, are mutually independent; see [12, Theorem 2.1] and Theorem 5.1. Consistent estimation of $T$ using the sample moment estimator $\widehat{T}_{S}$ requires a sample size satisfying
$N\gg\prod_{k=1}^{p}(r_{2}(\Sigma^{(k)})+\log N)^{1/2}\quad\text{or}\quad N\gg\prod_{k=1}^{p}(r_{\max}(\Sigma^{(k)})+\log N)^{1/2}$
under the operator norm and the entrywise maximum norm, respectively. In contrast, Isserlis’s estimator $\widehat{T}_{I}$ only requires sample size
$N\gg\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})\quad\text{or}\quad N\gg\max_{1\leq k\leq p}r_{\max}(\Sigma^{(k)})$
for consistency under the respective norms, leading to a significant reduction in sample complexity. ∎
Proof of Theorem 3.4.
First, for the sample moment estimator $\widehat{T}_{S}$, the upper bound on its deviation from $T$ under the operator norm follows directly from [12, Theorem 2.1], while the corresponding bound under the entrywise maximum norm is given by Theorem 5.1 in this paper. It remains to analyze Isserlis’s estimator $\widehat{T}_{I}$.
We apply the upper bound in Proposition 3.6, which yields
$\|\widehat{T}_{I}-T\|\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}(p-1)!!\cdot\frac{p}{2}\cdot\varepsilon_{*}(1+\varepsilon_{*})^{p/2-1},$
where
$\varepsilon_{*}:=\max_{j\neq k}\frac{\|\widehat{\Sigma}^{(j,k)}-\Sigma^{(j,k)}\|}{\big{(}\|\Sigma^{(j,j)}\|\|\Sigma^{(k,k)}\|\big{)}^{1/2}}.$
To control $\varepsilon_{*}$, we use the bound from [12, Remark 2.1 and page 21] for the sample cross-covariance. For $j,k\in\{1,\ldots,p\}$ and any $u\geq 0$, it holds with probability at least $1-\exp(-u^{2})$ that
$\frac{\|\widehat{\Sigma}^{(j,k)}-\Sigma^{(j,k)}\|}{\big{(}\|\Sigma^{(j,j)}\|\|\Sigma^{(k,k)}\|\big{)}^{1/2}}\lesssim u\bigg{(}\frac{r_{2}(\Sigma^{(j)})+r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+u^{2}\frac{(r_{2}(\Sigma^{(j)})r_{2}(\Sigma^{(k)}))^{1/2}}{N}.$
Applying a union bound over all $j\neq k$ yields, with probability at least $1-p^{2}\exp(-u^{2})$,
$\varepsilon_{*}$ $=\max_{j\neq k}\frac{\|\widehat{\Sigma}^{(j,k)}-\Sigma^{(j,k)}\|}{\big{(}\|\Sigma^{(j,j)}\|\|\Sigma^{(k,k)}\|\big{)}^{1/2}}$
$\lesssim\max_{j\neq k}\mathopen{}\mathclose{{}\left\{u\bigg{(}\frac{r_{2}(\Sigma^{(j)})+r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+u^{2}\frac{(r_{2}(\Sigma^{(j)})r_{2}(\Sigma^{(k)}))^{1/2}}{N}}\right\}$
$\lesssim u\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+u^{2}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}.$
Substituting this bound into the inequality for $\|\widehat{T}_{I}-T\|$, we obtain, with probability at least $1-p^{2}\exp(-u^{2})$:
$\|\widehat{T}_{I}-T\|$ $\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}(p-1)!!\cdot\frac{p}{2}\cdot\varepsilon_{*}(1+\varepsilon_{*})^{p/2-1}$
$\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}(\varepsilon_{*}+\varepsilon_{*}^{p/2})$
$\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}\bigg{(}u\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+u^{p}\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{p/2}\bigg{)}.$
Integrating the tail bound yields the following expectation bound:
$\mathbb{E}\|\widehat{T}_{I}-T\|$ $\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}\bigg{(}\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{p/2}\bigg{)}$
$\asymp_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|^{1/2}\bigg{)}\bigg{(}\frac{\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)})}{N}\bigg{)}^{1/2},\quad\text{ if }\ N\geq\max_{1\leq k\leq p}r_{2}(\Sigma^{(k)}).$
An analogous argument yields the corresponding bound under the entrywise maximum norm. This completes the proof. ∎
The following proposition is used in the proof of Theorem 3.4, and its proof is deferred to Section 4.
Proposition 3.6.
Let $p$ be an even integer, and let $X=(X^{(1)},\ldots,X^{(p)})$ and $Y=(Y^{(1)},\ldots,Y^{(p)})$ be zero-mean Gaussian random vectors in $\mathbb{R}^{d}$, where each $X^{(k)},Y^{(k)}\in\mathbb{R}^{d_{k}}$ and $d=\sum_{k=1}^{p}d_{k}$. For $j,k\in\{1,\ldots,p\}$, denote the cross-covariance blocks by
$\Sigma^{(j,k)}_{X}=\mathbb{E}X^{(j)}\otimes X^{(k)},\qquad\Sigma^{(j,k)}_{Y}=\mathbb{E}Y^{(j)}\otimes Y^{(k)}.$
Let
$T_{X}:=\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)},\qquad T_{Y}:=\mathbb{E}Y^{(1)}\otimes\cdots\otimes Y^{(p)},$
denote the moment tensors formed by taking tensor products over the blocks of $X$ and $Y$, respectively. Then,
$\|T_{X}-T_{Y}\|\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k,k)}_{Y}\|^{1/2}\bigg{)}(p-1)!!\cdot\frac{p}{2}\cdot\varepsilon_{*}(1+\varepsilon_{*})^{p/2-1},$
where
$\varepsilon_{*}:=\max_{j\neq k}\frac{\|\Sigma^{(j,k)}_{X}-\Sigma^{(j,k)}_{Y}\|}{\big{(}\|\Sigma^{(j,j)}_{Y}\|\|\Sigma^{(k,k)}_{Y}\|\big{)}^{1/2}}.$
For the entrywise maximum norm, it similarly holds that
$\|T_{X}-T_{Y}\|_{\max}\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k,k)}_{Y}\|_{\max}^{1/2}\bigg{)}(p-1)!!\cdot\frac{p}{2}\cdot\bar{\varepsilon}_{*}(1+\bar{\varepsilon}_{*})^{p/2-1},$
where
$\bar{\varepsilon}_{*}:=\max_{j\neq k}\frac{\|\Sigma^{(j,k)}_{X}-\Sigma^{(j,k)}_{Y}\|_{\max}}{\big{(}\|\Sigma^{(j,j)}_{Y}\|_{\max}\|\Sigma^{(k,k)}_{Y}\|_{\max}\big{)}^{1/2}}.$
Corollary 3.7.
Let $X\sim\mathcal{N}(0,\Sigma_{X})$ and $Y\sim\mathcal{N}(0,\Sigma_{Y})$, and let $T_{X}:=\mathbb{E}X^{\otimes p},T_{Y}:=\mathbb{E}Y^{\otimes p}$. Then,
$\frac{\|T_{X}-T_{Y}\|}{\|T_{Y}\|}\leq\frac{p}{2}\cdot\frac{\|\Sigma_{X}-\Sigma_{Y}\|}{\|\Sigma_{Y}\|}\mathopen{}\mathclose{{}\left(1+\frac{\|\Sigma_{X}-\Sigma_{Y}\|}{\|\Sigma_{Y}\|}}\right)^{p/2-1}.$
Similarly, under the entrywise maximum norm,
$\frac{\|T_{X}-T_{Y}\|_{\max}}{\|T_{Y}\|_{\max}}\leq\frac{p}{2}\cdot\frac{\|\Sigma_{X}-\Sigma_{Y}\|_{\max}}{\|\Sigma_{Y}\|_{\max}}\mathopen{}\mathclose{{}\left(1+\frac{\|\Sigma_{X}-\Sigma_{Y}\|_{\max}}{\|\Sigma_{Y}\|_{\max}}}\right)^{p/2-1}.$
Proof of Corollary 3.7.
Apply Proposition 3.6 with $X^{(1)}=\cdots=X^{(p)}=:X$ and $Y^{(1)}=\cdots=Y^{(p)}=:Y$, so that $\Sigma_{X}^{(j,k)}=:\Sigma_{X}$ and $\Sigma_{Y}^{(j,k)}=:\Sigma_{Y}$ for all $j,k\in\{1,\ldots,p\}$. The result then follows from computing the norm of the reference tensor:
$\|T_{Y}\|=\|\mathbb{E}Y^{\otimes p}\|=\sup_{v\in\mathcal{S}^{d}}\mathbb{E}\langle Y,v\rangle^{p}=(p-1)!!\cdot\|\Sigma_{Y}\|^{p/2},$
where the last equality uses the moment formula for centered Gaussian variables. The bound under the entrywise maximum norm follows analogously. ∎
4 Proof of Proposition 3.6
Proof of Proposition 3.6.
For arbitrary vectors $v_{1},\ldots,v_{p}\in\mathbb{R}^{d}$, we have
$\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle=\mathopen{}\mathclose{{}\left\langle\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}-\mathbb{E}Y^{(1)}\otimes\cdots\otimes Y^{(p)},v_{1}\otimes\cdots\otimes v_{p}}\right\rangle$
$=\mathbb{E}\prod_{k=1}^{p}\langle X^{(k)},v_{k}\rangle-\mathbb{E}\prod_{k=1}^{p}\langle Y^{(k)},v_{k}\rangle$
[…]
$=\sum_{\pi\in\Pi_{p}^{2}}\bigg{(}\prod_{(j,k)\in\pi}\langle\Sigma^{(j,k)}_{X}v_{j},v_{k}\rangle-\prod_{(j,k)\in\pi}\langle\Sigma^{(j,k)}_{Y}v_{j},v_{k}\rangle\bigg{)}.$
Here, (i) follows from Isserlis’s theorem, and $\Pi_{p}^{2}$ denotes the set of pairwise partitions of $\{1,\ldots,p\}$. Denoting by $\pi=(\pi(1),\ldots,\pi(p))$ a fixed ordering of the indices in the pairing $\pi$, and applying the telescoping identity
$a_{1}\cdots a_{p/2}-b_{1}\cdots b_{p/2}=\sum_{\ell=1}^{p/2}a_{1}\cdots a_{\ell-1}\mathopen{}\mathclose{{}\left(a_{\ell}-b_{\ell}}\right)b_{\ell+1}\cdots b_{p/2},$
we obtain
$\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle=\sum_{\pi\in\Pi_{p}^{2}}\bigg{(}\prod_{(j,k)\in\pi}\langle\Sigma^{(j,k)}_{X}v_{j},v_{k}\rangle-\prod_{(j,k)\in\pi}\langle\Sigma^{(j,k)}_{Y}v_{j},v_{k}\rangle\bigg{)}$
$=\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}$
×$\times\!\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\bigg{(}\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}.$
To bound the operator norm, we take the supremum over $v_{1},\ldots,v_{p}\in\mathcal{S}^{d}$,
$\|T_{X}-T_{Y}\|=\sup_{v_{1},\ldots,v_{p}\in\mathcal{S}^{d}}|\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle|$
$=\sup_{v_{1},\ldots,v_{p}\in\mathcal{S}^{d}}\bigg{|}\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}$
×$\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\times\bigg{(}\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\bigg{)}\bigg{|}$
$\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg{[}\sup_{v_{1},\ldots,v_{p}\in\mathcal{S}^{d}}\bigg{|}\bigg{(}\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}$
×$\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\times\bigg{(}\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\bigg{)}\bigg{|}\Bigg{]}$
$=\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{X}\|\bigg{)}\|\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y}\|\bigg{(}\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}\|\bigg{)}.$
For $j,k\in\{1,\ldots,p\}$, we introduce the normalized deviation $\varepsilon^{(j,k)}:=\frac{\|\Sigma^{(j,k)}_{X}-\Sigma^{(j,k)}_{Y}\|}{\big{(}\|\Sigma^{(j,j)}_{Y}\|\|\Sigma^{(k,k)}_{Y}\|\big{)}^{1/2}}$ and define $\varepsilon_{*}:=\max_{j\neq k}\varepsilon^{(j,k)}$. Using the inequality $\|\Sigma^{(j,k)}_{Y}\|\leq\|\Sigma^{(j,j)}_{Y}\|^{1/2}\|\Sigma^{(k,k)}_{Y}\|^{1/2}$, we obtain
$\|T_{X}-T_{Y}\|$
$\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{X}\|\bigg{)}\|\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y}\|\bigg{(}\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}\|\bigg{)}$
$\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg{[}\bigg{(}\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2s),\pi(2s))}_{Y}\|^{1/2}\big{(}1+\varepsilon^{(\pi(2s-1),\pi(2s))}\big{)}\bigg{)}$
×$\hskip 28.45274pt\times\bigg{(}\|\Sigma^{(\pi(2\ell-1),\pi(2\ell-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2\ell),\pi(2\ell))}_{Y}\|^{1/2}\varepsilon^{(\pi(2\ell-1),\pi(2\ell))}\bigg{)}$
×$\hskip 28.45274pt\times\bigg{(}\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2s),\pi(2s))}_{Y}\|^{1/2}\bigg{)}\Bigg{]}$
$=\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k,k)}_{Y}\|^{1/2}\bigg{)}\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\varepsilon^{(\pi(2\ell-1),\pi(2\ell))}\prod_{s=1}^{\ell-1}\big{(}1+\varepsilon^{(\pi(2s-1),\pi(2s))}\big{)}.$
Using the identity
$\sum_{\ell=1}^{p/2}a_{\ell}\prod_{s=1}^{\ell-1}(1+a_{s})=\prod_{\ell=1}^{p/2}(1+a_{\ell})-1,$
we conclude
$\|T_{X}-T_{Y}\|$ $\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k,k)}_{Y}\|^{1/2}\bigg{)}\sum_{\pi\in\Pi_{p}^{2}}\bigg{(}\prod_{\ell=1}^{p/2}\big{(}1+\varepsilon^{(\pi(2\ell-1),\pi(2\ell))}\big{)}-1\bigg{)}$
$\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k,k)}_{Y}\|^{1/2}\bigg{)}(p-1)!!\mathopen{}\mathclose{{}\left((1+\varepsilon_{*})^{p/2}-1}\right)$
$\leq\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k,k)}_{Y}\|^{1/2}\bigg{)}(p-1)!!\cdot\frac{p}{2}\cdot\varepsilon_{*}(1+\varepsilon_{*})^{p/2-1},$
where we used the inequality $(1+\varepsilon_{*})^{p/2}-1\leq\frac{p}{2}\cdot\varepsilon_{*}(1+\varepsilon_{*})^{p/2-1}$.
The bound under the entrywise maximum norm follows from an analogous argument, replacing the supremum over the unit sphere $\mathcal{S}^{d}$ with the supremum over the standard basis =$\mathcal{E}=\{e_{i}\}_{i=1}^{d}$. This completes the proof. ∎
5 Entrywise maximum norm bound for sample moment tensor
Theorem 5.1.
For any integer $p\geq 2$ and $1\leq k\leq p$, let $X^{(k)},X_{1}^{(k)},\ldots,X_{N}^{(k)}$ be i.i.d. zero-mean Gaussian random vectors in $\mathbb{R}^{d_{k}}$ with covariance matrix $\Sigma^{(k)}$. Then,
$\mathbb{E}\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}-\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}\bigg{\|}_{\max}\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|_{\max}^{1/2}\bigg{)}\mathscr{E}_{N}\big{(}(\Sigma^{(k)})_{k=1}^{p}\big{)},$
where
$\mathscr{E}_{N}\big{(}(\Sigma^{(k)})_{k=1}^{p}\big{)}:=\bigg{(}\frac{\sum_{k=1}^{p}r_{\max}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+\frac{1}{N}\prod_{k=1}^{p}\Big{(}r_{\max}(\Sigma^{(k)})+\log N\Big{)}^{1/2}.$
Moreover, the upper bound is sharp in the following two cases:
  1. 1.
    Independent Components. If $X^{(1)},\ldots,X^{(p)},(X^{(1)}_{i})_{i=1}^{N},\ldots,(X^{(p)}_{i})_{i=1}^{N}$ are mutually independent, then
    $\mathbb{E}\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}-\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}\bigg{\|}_{\max}\asymp_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|_{\max}^{1/2}\bigg{)}\mathscr{E}_{N}\big{(}(\Sigma^{(k)})_{k=1}^{p}\big{)}.$
  2. 2.
    Identical Components. If $X^{(1)}=\cdots=X^{(p)}=X$ and $X_{i}^{(1)}=\cdots=X_{i}^{(p)}=X_{i}\,$ for all $1\leq i\leq N$, $\Sigma^{(1)}=\cdots=\Sigma^{(p)}=\Sigma,$ then
    $\mathbb{E}\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}-\mathbb{E}X^{\otimes p}\bigg{\|}_{\max}$ $\asymp_{p}\|\Sigma\|_{\max}^{p/2}\bigg{(}\sqrt{\frac{r_{\max}(\Sigma)}{N}}+\frac{(r_{\max}(\Sigma)+\log N)^{p/2}}{N}\bigg{)}.$
    Furthermore, since $(\log N)^{p/2}/N\lesssim_{p}1/\sqrt{N}$, this bound simplifies to
    $\mathbb{E}\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}-\mathbb{E}X^{\otimes p}\bigg{\|}_{\max}\asymp_{p}\|\Sigma\|_{\max}^{p/2}\bigg{(}\sqrt{\frac{r_{\max}(\Sigma)}{N}}+\frac{r_{\max}(\Sigma)^{p/2}}{N}\bigg{)}.$
Remark 5.2.
The upper bound in Theorem 5.1 holds without requiring independence between the sequences $(X^{(k)}_{i})_{i=1}^{N}$ and $(X^{(k^{\prime})}_{i})_{i=1}^{N}$ for $k\neq k^{\prime}$; that is, no assumptions are made on the correlation structure across components. Moreover, the upper bound in Theorem 5.1 extends directly to sub-Gaussian settings. Dimension-dependent counterparts can also be derived by applying standard $\varepsilon$-net arguments in conjunction with the $\alpha$-sub-exponential concentration inequality of [14]. ∎
Proof of Theorem 5.1.
Upper bound By the definition of the entrywise maximum norm, we have
$\mathbb{E}\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}-\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}\bigg{\|}_{\max}$
$=\mathbb{E}\max_{v_{1},\dots,v_{p}\in\mathcal{E}}\bigg{|}\frac{1}{N}\sum_{i=1}^{N}\prod_{k=1}^{p}\langle X^{(k)}_{i},v_{k}\rangle-\mathbb{E}\prod_{k=1}^{p}\langle X^{(k)},v_{k}\rangle\bigg{|},$
where $\mathcal{E}$ denotes the standard basis of $\mathbb{R}^{d}$. Let $\bar{\mathcal{E}}:=\mathcal{E}\cup-\mathcal{E}$, and define $\mathcal{F}^{(1)}=\cdots=\mathcal{F}^{(p)}:=\{\langle\cdot,v\rangle:v\in\bar{\mathcal{E}}\}$. Then the maximum can be upper bounded by
$\mathbb{E}\sup_{f^{(k)}\in\mathcal{F}^{(k)},1\leq k\leq p}\bigg{|}\frac{1}{N}\sum_{i=1}^{N}\prod_{k=1}^{p}f^{(k)}(X^{(k)}_{i})-\mathbb{E}\prod_{k=1}^{p}f^{(k)}(X^{(k)})\bigg{|}.$
Applying [12, Theorem 2.2], we obtain
$\mathbb{E}\!\sup_{f^{(k)}\in\mathcal{F}^{(k)},1\leq k\leq p}\bigg{|}\frac{1}{N}\sum_{i=1}^{N}\prod_{k=1}^{p}f^{(k)}(X^{(k)}_{i})-\mathbb{E}\prod_{k=1}^{p}f^{(k)}(X^{(k)})\bigg{|}$
$\lesssim_{p}\bigg{(}\prod_{k=1}^{p}d_{\psi_{2}}(\mathcal{F}^{(k)})\bigg{)}\bigg{(}\frac{\sum_{k=1}^{p}\bar{\gamma}(\mathcal{F}^{(k)},\psi_{2})}{\sqrt{N}}+\frac{\prod_{k=1}^{p}\big{(}\bar{\gamma}(\mathcal{F}^{(k)},\psi_{2})+(\log N)^{1/2}\big{)}}{N}\bigg{)},$ (5.1)
where $\bar{\gamma}(\mathcal{F}^{(k)},\psi_{2}):=\gamma(\mathcal{F}^{(k)},\psi_{2})/d_{\psi_{2}}(\mathcal{F}^{(k)})$. Here, $\gamma(\mathcal{F},\psi_{2})$ denotes Talagrand’s generic chaining complexity of the function class $\mathcal{F}$ [26, Definition 2.7.3], $d_{\psi_{2}}(\mathcal{F}):=\sup_{f\in\mathcal{F}}\|f\|_{\psi_{2}}$, and $\psi_{2}$ refers to the Orlicz norm with Orlicz function $\psi(x)=e^{x^{2}}-1,$ see e.g. [29, Definition 2.5.6].
Since $X^{(k)}$ is Gaussian with covariance $\Sigma^{(k)}$, the $\psi_{2}$-norm of linear functionals is equivalent to the $L_{2}$-norm. Hence,
$d_{\psi_{2}}(\mathcal{F}^{(k)})=\sup_{f^{(k)}\in\mathcal{F}^{(k)}}\|f^{(k)}\|_{\psi_{2}}\asymp\sup_{f^{(k)}\in\mathcal{F}^{(k)}}\|f^{(k)}\|_{L_{2}}=\sup_{v\in\bar{\mathcal{E}}}\big{(}\mathbb{E}\langle X^{(k)},v\rangle^{2}\big{)}^{1/2}=\|\Sigma^{(k)}\|_{\max}^{1/2}.$
For the generic chaining term, note that the canonical metric on $\bar{\mathcal{E}}$ is given by
$d_{X^{(k)}}(u,v):=\big{(}\mathbb{E}(\langle X^{(k)},u\rangle-\langle X^{(k)},v\rangle)^{2}\big{)}^{1/2}=\langle u-v,\Sigma^{(k)}(u-v)\rangle^{1/2}=\|\langle\cdot,u\rangle-\langle\cdot,v\rangle\|_{L_{2}(\mu^{(k)})},$
where $\mu^{(k)}$ is the law of $X^{(k)}$. By Talagrand’s majorizing measure theorem [26, Theorem 2.10.1],
$\gamma(\mathcal{F}^{(k)},\psi_{2})\asymp\gamma(\mathcal{F}^{(k)},L_{2})=\gamma(\bar{\mathcal{E}},d_{X^{(k)}})\asymp\mathbb{E}\sup_{u\in\bar{\mathcal{E}}}\langle X^{(k)},u\rangle=\mathbb{E}\|X^{(k)}\|_{\infty}.$
We conclude that
$d_{\psi_{2}}(\mathcal{F}^{(k)})\asymp\|\Sigma^{(k)}\|_{\max}^{1/2},\quad\gamma(\mathcal{F}^{(k)},\psi_{2})\asymp\mathbb{E}\|X^{(k)}\|_{\infty},$
and therefore,
$\mathbb{E}\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}-\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}\bigg{\|}_{\max}\lesssim_{p}\bigg{(}\prod_{k=1}^{p}\|\Sigma^{(k)}\|_{\max}^{1/2}\bigg{)}\mathscr{E}_{N}\big{(}(\Sigma^{(k)})_{k=1}^{p}\big{)},$
where
$\mathscr{E}_{N}\big{(}(\Sigma^{(k)})_{k=1}^{p}\big{)}:=\bigg{(}\frac{\sum_{k=1}^{p}r_{\max}(\Sigma^{(k)})}{N}\bigg{)}^{1/2}+\frac{1}{N}\prod_{k=1}^{p}\Big{(}r_{\max}(\Sigma^{(k)})+\log N\Big{)}^{1/2}$
and $r_{\max}(\Sigma^{(k)}):=\mathopen{}\mathclose{{}\left(\mathbb{E}_{X^{(k)}\sim\mathcal{N}(0,\Sigma^{(k)})}\|X^{(k)}\|_{\infty}}\right)^{2}/\|\Sigma^{(k)}\|_{\max}$.
Lower bound The lower bounds follow from straightforward modifications of the argument in [12, Proposition 3.1] and [2, Proposition 3.1], respectively, by replacing the operator norm with the entrywise maximum norm. We omit the details for brevity. ∎
Acknowledgments
The work of DSA was partly funded by NSF CAREER DMS-2237628. The work of OAG was supported in part by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard.